OpenStudy (anonymous):

Solve the absolute value inequality: |5x + 3| (less than or equal to) 13 a. x -3.2 and x 2 b. x -3.2 or x 2 c. x -3.2 or x 2 d. x -3.2 or x 3.2

OpenStudy (anonymous):

\[|5x + 3| \le 13\] \[-13 \le 5x+3 \le 13\] \[-16 \le 5x \le 10\] \[-16/5 \le x \le 2\] so: \[-3.2 \le x \le 2\]

OpenStudy (anonymous):

It had this sign  so im really not sure what it ment?

OpenStudy (anonymous):

btw thanks for solving but i don't unerstand